OpenStudy (abbycross167):

Write the slope-intercept (y = mx + b) form of the equation from the information given. The line goes through the points (-1, -7) and (-8, -2) Slope = y2 - y1 =-1 - (-2) x2 - x1 -7 - (-8) = 2/1 @AbdullahM this is all I could figure out

OpenStudy (abdullahm):

@abbycross167 y2 - y1 would be -7 - (-2) = -7 + 2 = -5 x2 - x1 = -1 - (-8) = -1 + 8 = 7 So the slope is -5/7